OpenStudy (anonymous):

1/5 - ( ? ) = 3/40 ? please help.

OpenStudy (anonymous):

Gummage friend, (:

OpenStudy (anonymous):

(3/40)-(1/5)=? 1/5 = 8/40 so 3/40-8/40 =-5/40 -1/8

OpenStudy (anonymous):

so, whats the answer?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

as -1/8=-(?)